OpenStudy (anonymous):

solve the equation using quadratic formula x^2+10x+29=0 please show work

OpenStudy (saifoo.khan):

a = 1 b = 10 c = 29 Inserting these values in quadratic formula, \[\frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]\[\frac{-10 \pm \sqrt{10^2 - 4(1)(29)}}{2(1)}\]

OpenStudy (anonymous):

thanks!

OpenStudy (saifoo.khan):

Welcome.